Factor each of the trinomials, If the question does not factor write none
INTERMEDIATE ALGEBRA
X2-2XY-15Y2=
5X2+25X+30=
X2-XY-12Y2=
X2-7X+12=

Answer by stanbon(75887)

X2-2XY-15Y2=
(x + 3y)(x - 5y)
---------------------
-
5X2+25X+30
= 5(x^2 + 5x + 6)
=5(x+3)(x+2)
-------------------------

X2-XY-12Y2
= (x-4y)(x+3y)
------------------------

X2-7X+12
=(x-3)(x-4)
====================
Cheers,
Stan H.